### v2

# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

Abstract.
When allowing concurrent actions in Markov Decision Processes, whose state and action spaces grow exponentially in the number of objects, computing a policy becomes highly inefficient, as it requires enumerating the joint of the two spaces. For the case of indistinguishable objects, we present a first-order representation to tackle the exponential blow-up in the action and state spaces. We propose Foreplan, an efficient relational forward planner, which uses the first-order representation allowing to compute policies in space and time polynomially in the number of objects. Thus, Foreplan significantly increases the number of planning problems solvable, in an exact manner, in reasonable time, which we underscore with a theoretical analysis. To speed up the computations even further, we also introduce an approximate version of Foreplan, including guarantees on the error. Further, we provide an empirical evaluation of both Foreplan versions, demonstrating a speedup of several orders of magnitude. For the approximate version of Foreplan, we also empirically show that the induced error is often negligible.
Key words and phrases:
Planning, Markov Decision Process, Lifting, Probabilistic Graphical Models
1. Introduction
To compute a policy for a Markov decision processes (MDPs), combinations of the state and action spaces have to be enumerated to find the best possible action for each state. With increasing number of (indistinguishable) objects, the state space grows exponentially, leading to drastically more combinations to be enumerated. In case the actions also depend on the states or objects, the action space grows as well, leading to even more combinations for enumeration. Further, if one allows concurrent actions, all possible combinations of the actions, yielding an exponentially-sized action space, of the already enlarged action space have to be used during the enumeration. Therefore, given the number of possibilities that have to be enumerated, concurrency is rarely modelled. However, consider the following example: A small town is haunted by an epidemic. To fight the epidemic, the town’s mayor can impose travel bans on the town’s citizens. Certainly, the mayor can confine all citizens to their homes, stopping the epidemic. However, the citizens’ overall welfare is important as well. Therefore, the mayor is interested in the best decision of imposing travel bans (concurrent actions) w.r.t. confining the epidemic, while keeping the citizens’ welfare above a certain threshold. In addition, there can be circumstances where the policy has to be adapted, e.g., if a wave of infections is on the rise, threatening to infect the majority of the population. Computing this problem on a propositional level, with every citizen explicitly represented, like approaches such as MDPs do, blows up the state and action space, as it requires exponential enumeration of all subsets of the population. However, there are groups of citizens behaving identically w.r.t. getting sick (and well again) as well as regarding their welfare if a travel ban is imposed, i.e., these citizens are indistinguishable for the mayor. Within these groups, it does not matter on which exact citizen the mayor imposes a travel ban on, only on how many she imposes a travel ban on. Additionally, many computations over subsets of indistinguishable citizens are redundant. Thus, we propose to drastically reduce the search and action space by using a first-order representation, grouping indistinguishable citizens, and use this representation to plan group-wise by reasoning about a representative and then projecting the result to the whole group.
Contribution
First, to be able to represent numerous indistinguishable objects and actions, we use probabilistic relational models as a representation in (factored) MDPs (fMDPs), which yields relational factored MDPs (rfMDPs). Second, as our key contribution, we propose Foreplan to carry out exact planning in rfMDPs. Foreplan uses our first-order representation to keep the action and state space only polynomial in the number of objects, which we prove by using our so-called relational cost graph. Foreplan remains exponential in the number of cliques $c$ and size $w$ of the largest clique in the relational cost graph, but both parameters are mostly small and independent of the domain sizes. Therefore, Foreplan is an efficient planner w.r.t the number of objects. We show a speedup of at least four orders of magnitude is achievable. Last, following the approximation ideas of approx_factored_mdps, we can reduce the runtime even further. We propose Approximate Foreplan, whose runtime is polynomial in $c$. For Approximate Foreplan, we show a speedup of at least four orders of magnitude.
Related Work
mdp introduces MDPs, which factored_mdp extend to fMDPs by factorizing the transition function. Factorizing also the value function, approx_factored_mdps provide two approximate algorithms for solving planning in fMDPs. dean2013model cluster the state space of fMDPs to reduce the state space even further. givan2003equivalence group equivalent states based on the notion of bisimulation. Both approaches lack the ability to handle concurrent actions efficiently. MDPs can be generalized to partially observable MDPs, in which the agent is uncertain in which state the environment currently is kaelbling1998planning. sanner2010symbolic add the first-order perspective to partially observable MDPs, focusing on the observations and without concurrent actions. bernstein2002complexity extend partially observable MDPs to have a set of decentralized agents. braun2022lifting work with groups of indistinguishable agents. This is similar to our approach, in which we handle sets of indistinguishable state variables. However, they do not provide an algorithm.
The idea of lifting is to carry out computations over representatives for groups of indistinguishable random variables lve; kersting2012lifted; raedt2016statistical. There are online decision making approaches adding action and utilitity nodes to this representation apsel2011extended; 10.1007/978-3-030-12738-1_10; 10.1007/978-3-030-18305-9_33, here, we focus on offline planning. To carry out even more lifted computations, crv extends lifted probabilistic inference by a generalized counting framework, which we extend later on. Using a first-order representation for states, actions and objects, boutilier2001symbolic exploit the relational structure for planning using MDPs, still without concurrent actions. To specify factored transition models in first-order MDPs, sanner2007approximate introduce factored first-order MDPs, which uses a backward search and provide only an approximate algorithm. In contrast, we propose an exact algorithm by applying lifting. Later, we also introduce an approximate version. Moreover, we prove on which models our algorithm runs efficiently. For a survey on planning using first-order representations, we refer to correa2024lifted.
Structure
The remainder of this paper is structured as follows: First, we present preliminaries for (factored) MDPs, including lifting. Then, we introduce rfMDPs to group indistinguishable objects. Afterwards, we propose Foreplan, which exploits these indistinguishable objects using a compact state representation, to efficiently support decision making with concurrent actions. Further, we provide a theoretical analysis and empirical evaluation of Foreplan.
2. Preliminaries
In this section, we lay the foundation for rfMDPs. We first recap (f)MDPs, which model probabilistic state changes, when an agent performs an action. Furthermore, the states have a reward assigned, and the task is to compute the optimal policy, i.e., which action to perform in which state, for the agent w.r.t. rewards. Second, we recap lifting in probabilistic graphical models.
2.1. (Factored) Markov Decision Processes
In this subsection, we first define MDPs and specialize them to fMDPs by factoring the transition function $T$.
[…]
A Markov Decision Process is a tuple $(\mathbf{S},\mathbf{A},T,R)$ with a set of states $\mathbf{S}$, a set of actions $\mathbf{A}$, a reward function $R:\mathbf{S}\mapsto\mathbb{R}$ and a transition function $T:\mathbf{S}\times\mathbf{A}\times\mathbf{S}\to[0,1]$. The value $T(s,a,s^{\prime})$ is the probability $P(s^{\prime}\mid s,a)$ of transitioning from state $s\in\mathbf{S}$ to state $s^{\prime}\in\mathbf{S}$ if action $a\in\mathbf{A}$ is performed.
The rewards are additive, possibly discounted by a factor $\gamma\in[0,1)$. An MDP is fully observable and has the first order Markov property, i.e., the probability of the next state only depends on the current state and action. Let us have a look at a simple, (incomplete) example of an MDP.
[…]
Suppose we have the states healthy and sick. Our agent has two possible actions: Travelling or staying at home. When the agent is in state sick and travels, she stays sick with a probability of $0.9$. The agent obtains a reward of 1-1 if sick and $1$ if healthy.
Planning in MDPs refers to calculating an optimal policy for the agent, which is a mapping from each state to an action to perform. To compute such a policy, we first define the utility of a state:
[…]
The utility of a state $s$ is given by
$U(s)=R(s)+\gamma\max_{a\in\mathbf{A}}\sum_{s^{\prime}\in\mathbf{S}}P(s^{\prime}\mid s,a)\cdot U(s^{\prime}).$ (1)
To find the utility of a state algorithmically, we find a value function $V$ satisfying the Bellman equation. The value function induces a policy by selecting the action that yields the maximum expected value. For computing a value function, we can use a linear programming formulation mdp_lp_orig; mdp_lp:
[…]
where the coefficients $\alpha(s)$ are arbitrary positive numbers, e.g., an equal distribution over all states mdp_lp. Planning in MDPs can be solved in polynomial time w.r.t to the state space size papadimitriou1987complexity. But what if the state space becomes very large, e.g., exponential in the number of objects? For retaining an efficient transition model, fMDPs make use of state variables for the objects. The state space is then spanned by the state variables and is exponential in the number of state variables. For simplicity, all state variables are Boolean, but can be easily extended to non-Boolean.
[…]
A factored MDP is a tuple $(\mathbf{S}$, $\mathbf{A}$, $T$, $R)$, where the state space is the cross product $\mathbf{S}=\{0,1\}^{m}$ of the sub-spaces corresponding´to the state variables $S_{1},\dots,S_{m}$ with Boolean range. The transition function $T$ is, for each action, factored:
$P(S^{\prime}\mid S,a)=\prod_{i=1}^{m}P(S_{i}^{\prime}\mid Pa(S_{i}^{\prime}),a),$ (3)
where $Pa(S_{i}^{\prime})\subseteq\mathbf{S}$ denotes the set of parents of $S_{i}^{\prime}$ in the previous state, $S$ the old state, $S^{\prime}$ the new state and $S_{i}^{\prime}$ the $i$-th state variables in the respective state. For a given state $s\in\mathbf{S}$, we denote with $s_{i}$ the assignment of state variable $S_{i}$ in state $s$.
When having numerous objects, the state space explodes. But when these objects are indistinguishable, we can use parameterized graphical models to encode the state space and transition model compactly.
2.2. Parameterized Graphical Models
Having indistinguishable random variables leads to the redundant computations. We can tackle redundant computations by parameterizing our probabilistic model and grouping indistinguishable variables, so that inference in the probabilistic model becomes tractable w.r.t. domain sizes by using representatives during calculations niepert2014tractability. We first define parameterized random variables (PRVs) to group same behaving variables:
[…]
Let $\mathbf{W}$ be a set of random variable names, $\mathbf{L}$ a set of logical variable (logvar) names, and $\mathbf{D}$ a set of constants (universe). All sets are finite. Each logvar $L$ has a domain $\mathcal{D}(L)\subseteq\mathbf{D}$. A constraint $C$ is a tuple $(\mathcal{X},C_{\mathcal{X}})$ of a sequence of logvars $\mathcal{X}=(X_{1},\dots,X_{n})$ and a set $C_{\mathcal{X}}\subseteq\times_{i=1}^{n}\mathcal{D}(X_{i})$. The symbol $\top$ for $C$ marks that no restrictions apply, i.e., $C_{\mathcal{X}}=\times_{i=1}^{n}\mathcal{D}(X_{i})$. A PRV $B(L_{1},\dots,L_{n}),n\geq 0,$ is a syntactical construct of a random variable $B\in\mathbf{W}$ possibly combined with logvars $L_{1},\dots,L_{n}\in\mathbf{L}$. If $n=0$, the PRV is parameterless and constitutes a propositional random variable (RV). The term $\mathcal{R}(B)$ denotes the possible values (range) of a PRV $B$. An event $B=b$ denotes the occurrence of PRV $B$ with range value $b\in\mathcal{R}(B)$.
We create a PRV for the health status of each person in the town:
[…]
Let $\mathbf{W}=\{Sick,Epidemic\}$, $\mathbf{L}=\{M\}$, $\mathcal{D}(M)=\mathbf{D}=\{a,b,c,d,e,f,g,h\}$ with Boolean-valued PRVs $Sick(M)$ and $Epidemic$.
EpidemicEpidemic is a parameterless PRV. To set PRVs into relation, we use parameterized factors:
[…]
Let $\Phi$ be a set of factor names. We denote a parameterized factor (parfactor) $g$ by $\phi(\mathcal{B})_{|C}$ with $\mathcal{B}=(B_{1},\dots,B_{n})$ a sequence of PRVs, $\phi:\times_{i=1}^{n}\mathcal{R}(B_{i})\to\mathbb{R}^{+}$ a potential function with name $\phi\in\Phi$, and $C$ a constraint on the logvars of $\mathcal{B}$. A set of parfactors forms a model $G:=\{g_{i}\}_{i=1}^{n}$. With $Z$ as normalizing constant, $G$ represents the full joint distribution $P_{G}=\frac{1}{Z}\prod_{f\in gr(G)}f$, with $gr(G)$ referring to the groundings of $G$ w.r.t. given constraints. A grounding is the instantiation of each parfactor with an allowed constant.
Continuing Example 2.2, let us define the parfactor model:
[…]
Let $\forall m\in M:\phi(Sick(m),Epidemic)_{\top}$ be the parfactor with potential $\phi$ defining the probability to be sick for all persons $\mathbf{D}$, given there is an epidemic (or not). The grounded model then consists of the eight factors $\phi(Sick(a),Epidemic)$, $\phi(Sick(b)$, $Epidemic)$, …, $\phi(Sick(h),Epidemic)$ compared to one parfactor in the lifted model. Table 1 shows the potential function.
Table 1. Potential function giving the probability of a person being sick given whether there is an epidemic.
EpidemicEpidemic $Sick(M)$
0 0 0.9
0 1 0.1
1 0 0.3
1 1 0.7
Next, we integrate the parfactor into fMDPs to achieve a compact representation of the state and action spaces.
3. Relational Factored MDPs
We present the first-order representation for the state and action space as well as the transition function, which allows to efficiently represent indistinguishable objects as well as concurrent action, which then in turn can be exploited by an appropriate algorithm. To that end, we present Foreplan in the next section.
We define rfMDPs based on fMDPs, but include a parfactor model inside the state and action space and transition function. That is, the state variables can now contain PRVs, which are then used in the transition model and the reward function. Also, we have action PRVs, whose value is chosen by the agent. Basically, we keep the same semantics as in Definition 2.1 and only change the representation to exploit indistinguishable variables. In the following definition we use the term interpretation of a set for a truth-value assignment to each element in the set.
[…]
A relational factored MDP is a tuple $(\mathbf{D},\mathbf{L},\mathbf{X},\mathbf{A},G,\mathbf{R})$. The set $\mathbf{D}$ is a set of constants and the set $\mathbf{L}$ is a set of logvars over $\mathbf{D}$. The set $\mathbf{X}$ is a set of PRVs defined over $\mathbf{L}$. The set of possible interpretations $\mathcal{I}_{\mathbf{X}}$ for the groundings of the set $\mathbf{X}$ defines the state space, i.e., all possible assignments. The set $\mathbf{A}$ is a set of action PRVs. A parfactor model $G$ over $\mathbf{A}$ and $\mathbf{X}$ represents the transition function $T:\mathcal{I}_{\mathbf{X}}\times\mathcal{I}_{\mathbf{A}}\times\mathcal{I}_{\mathbf{X}}\to\mathbb{R}_{0}^{+}$, with the set $\mathcal{I}_{\mathbf{A}}$ of possible interpretations of the groundings of $\mathbf{A}$, and specifies the transition probability given an action and a previous state. The set $\mathbf{R}$ contains parameterized local reward functions $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\to\mathbb{R}$, defined over all PRVs $B_{i,j}$, which contribute to the local reward $R_{i}$. The reward function $R$ is decomposed as a sum over $\mathbf{R}$.
Let us generalize Example 2.1 to an arbitrary number of persons behaving in the same way as the agent in Example 2.1:
[…]
There is a set $\mathbf{D}$ of persons living in a small town, represented by the logvar $M$ with $\mathcal{D}(M)=\mathbf{D}$. Each person can be sick or healthy, leading to the PRV $Sick(M)$. The government gets a reward of $1$ for each healthy person and a reward of 1-1 for each sick person. To combat an epidemic, the government can impose travel bans on persons, resulting in the action $Restrict(M)$ to impose a travel ban on a subset of persons. Moreover, each person can travel or not, leading to the PRV $Travel(M)$. The government gets a reward of $2$ for each person travelling. The PRV $Epidemic$ is influenced by the number of people travelling and influences the healthiness of each person. Figure 1 shows the grounded transition model and Figure 2 shows the lifted transition model for this example.
[…]
Figure 1. Grounded transition model for Example 3. We assume that the population of the town consists of two citizens $a$ and $b$. For abbreviation, we use only the first letter(s) for each symbol. The factors $f_{1}$ and $\tilde{f}_{1}$ as well as $f_{2}$ and $\tilde{f}_{2}$ could be combined into a parfactor. Figure 2 shows the lifted version.
[…]
Grounded graphical model for the epidemic example: The factor f_1 links Travel(a) and Restrict(a) to Travel’(a), analogously tilde f_1 for b, f_2 links Sick(a) and Epidemic to Sick’(a), analogously tilde f_2 for b, and f_3 links Travel(a) and Travel(b) to Epidemic’.
[…]
Figure 2. Lifted representation of the transition model for Example 3. We abbreviate by using only the first letter(s) for each symbol.
[…]
Since an action can be applied to each person concurrently, the amount of possible actions is exponential due to the power set, i.e., all possible combinations for all persons. The action PRVs help Foreplan to reduce the exponential blow-up by exploiting the indistinguishability.
With Definition 3, we require the rewards to be represented as a sum of local reward functions, and not as some arbitrary function. Actually, this is too restrictive: It is sufficient that an operation running over same behaving variables has a hyperoperation repeating the operation multiple times. For a sum, the hyperoperation is multiplication. Also, multiple terms of such operations can be linked together in any way. In Appendix A.1, we give examples for hyperoperations. For simplicity, we use the definition as it is throughout the paper.
While Definition 3 defines the ground semantics, there are symmetries within the parfactor model, which we exploit with Foreplan in the next section. In the remainder of this section, we explain what we mean by action PRVs and by parameterized local reward functions.
3.1. Parameterizing Actions
In our epidemic example, the mayor, representing the government, can impose travel bans on all parts of the town’s population. We extend the action definition in this subsection to account for groups of objects. Having groups, we circumvent enumeration of all possible subsets and model the example action of imposing travel bans on a subset of the population efficiently.
[…]
An action PRV $A$ is a Boolean-valued PRV. A concrete action is a set of events, i.e., each grounding of $A$ receives an assignment $a\in\mathcal{R}(A)$.
Action PRVs allow for a more general action setting. When writing action, we refer to Definition 3.1, i.e., concurrent actions. In our example, the mayor can restrict multiple persons from travelling at once:
[…]
The action PRV $Restrict(M)$ models the possible travel bans on the population of the town. For a concrete action, the mayor has to specify which persons are restricted from travelling. When restricting the persons $a$, $b$ and $f$, the mayor specifies $Restrict(M)=true$ with constraint $(M,\{a,b,e\})$ and $Restrict(M)=false$ with constraint $(M,\{c,d,f,g,h\})$.
When the persons $a$ to $h$ are indistinguishable, it is irrelevant which permutation is used for the constraints, only the amounts are relevant, which we exploit in Foreplan.
As actions are now parameterized, we describe the impact of parameterization on the rewards.
3.2. Parameterized Local Reward Functions
The reward function in fMDPs maps from the (joint) state to the reward of the state. For evaluating the reward function, we thus have to construct the joint state and cannot exploit the factorization, breaking our aim of compact representation. To further use our compact representation, we introduce a decomposable reward function: For simplicity, we assume that the reward function is factored as $R=\sum_{i}R_{i}$, with local reward functions $R_{i}$ with scope restricted to a subset of the state variables. Other operations are possible too, as long as they have a hyperoperation, as already mentioned. As we have indistinguishable state variables, we reduce redundant computations by using representatives in the reward functions analogue to a parfactor, but using a sum instead of a product:
[…]
A local reward function $R_{i}:\bigtimes_{j}\mathcal{R}(B_{i,j})\to\mathbb{R}$ is defined over all PRVs $B_{i,j}$, which contribute to the local reward $R_{i}$. The semantics of a single parameterized local reward function $R_{i}$ is defined as the sum $\sum_{z}R_{i}(z)$ over the interpretations $z\in\bigtimes_{j}\mathcal{R}(B_{i,j})$ in the current state of all groundings of $R_{i}$.
In other words, a parameterized local reward function serves as a placeholder for the set of local reward functions, obtained by replacing all logvars by their possible instantiations. We illustrate parameterized reward functions in our epidemic example:
[…]
The parameterized local reward functions for Example 3 are $R_{1}(Sick(M))$, evaluating to 1-1 (11) for each person (not) being sick, and $R_{2}(Travel(M))$, evaluating to $2$ for each person travelling. If five persons are sick, three are not sick and four people are travelling, the total reward is 5+3+8=6-5+3+8=6.
With rfMDPs, we can efficiently represent the action and state space as well as the transition function for numerous indistinguishable objects. Further, we have a compact representation for concurrent actions. In the next section, we propose Foreplan, our main contribution, which uses our representation to solve planning in rfMDPs in time polynomial in the number of objects.
4. Foreplan: Efficient Forward Planning
In this section, we propose Foreplan, our exact forward relational planner for rfMDPs. The input to Foreplan is an rfMDP. The output is a value function, which induces a policy. Foreplan computes the value function by using a compact state representation for the rfMDP to exploit the indistinguishability of the objects and then running a linear program based on the state representation to calculate the value function.
The first-order representation in rfMDP contains dependencies between PRVs. For using the representation in computation, we first need to recognize the dependencies and find a suitable representation of them to compute the value function. We first describe how Foreplan recognizes the dependencies and identifies a suitable representation, Afterwards, we outline how Foreplan uses the identified representation to compute a value function.
4.1. Obtaining a Compact State Representation
Foreplan needs to encode the current state compactly to efficiently reason about indistinguishable variables. Thus, in this subsection, we describe how Foreplan treats indistinguishable objects for computations using rfMDPs. The representation in rfMDPs does not consider dependencies between PRVs. But, Foreplan can tame the information to store: Foreplan does not need to keep track of objects that could be differentiated by their history. Rather, with each action and new time step, the history is swept away and the objects remain indistinguishable because of the first-order Markov assumption. Thus, the basic idea is to count the objects for which the PRVs are true or false, which is inspired by Counting Random Variables (CRVs) crv. When evaluating a parfactor, the joint assignment to the PRVs is needed, rather than individual counts for each PRV. It is sufficient to count the number of occurrences of each possible truth-value assignment to the groundings of the input PRVs of a parfactor in a histogram. The input PRVs of a parfactor are the ones representing the current state. Counting only the input PRVs is sufficient, because all possible next states are iterated separately, which we describe in the next subsection. But some parfactors may be defined over PRVs defined over different logvars, like $f_{3}$ in Figure 2 is defined over $Travel(M)$ and $Epidemic$. Given the counts for $Travel(M)$ and $Epidemic$ separately is enough, since these PRVs do not depend on each other. But how does Foreplan recognize which PRVs depend on which PRVs? For answering this question, we propose the relational cost graph to identify the dependencies between PRVs and identify the groups that need to be counted jointly. Focusing only on the counts for the groups enables Foreplan to use a much simpler state space representation, namely the set of possible histograms, compared to the grounded state space representation, which is the Cartesian product over all state variables. We now describe in more detail how to count the assignments.
Counting the assignments of each PRV separately is insufficient, as PRVs can be defined over the same logvars and thus interfere with each other. However, the parfactors can be evaluated separately since, in Equation 2, we have the full current and next state available. Thus, it is sufficient to count PRVs together if they share a logvar and occur in a parfactor together. To obtain the representation and later on quantify its complexity, we define the relational cost graph:
[…]
The relational cost graph of an rfMDP has a vertex for each PRV in the current state. Two vertices are connected by an edge if and only if the PRVs associated with these two vertices share a logvar and occur together in a parfactor as input PRVs or a parameterized local reward function. We denote the number of (maximal) cliques by $c$ and the size of the largest clique by $w$.
We give a more formal definition of the relational cost graph in Appendix A.2. Let us take a look at the relational cost graph of Example 3. For a more complex example, we refer to Appendix A.3.
[…]
The relational cost graph for Example 3 consists of three isolated vertices corresponding to $Sick(M)$, $Travel(M)$ and $Epidemic$. The first two do not occur together in a parfactor or local reward function and the last does not share a logvar with one of the first two. Figure 3 shows the visual representation of the relational cost graph.
Si(M)Si(M)Tr(M)Tr(M)EpiEpi
Figure 3. Visual representation of the relational cost graph in Example 3. We abbreviate using the first letter(s).
[…]
The relational cost graph for the epidemic example consists of three isolated vertices corresponding to Sick(M), Travel(M) and Epidemic.
The key insight now is that (maximal) cliques in the relational cost graph correspond to sets of PRVs that Foreplan needs to count together as they interfere with each other. But when having more than one logvar in a PRV, basic CRVs crv only allow to count one of them. To use our representation in computations, we extend CRVs to allow for more than one logvar, extending the definition by crv:
[…]
A counting formula $\gamma=\#_{C}[B_{1},\dots,B_{k}]$ is defined over PRVs $B_{i}$ with a constraint $C=(\mathcal{L},C_{\mathcal{L}})$ over the logvars $\mathcal{L}$ of the PRVs $B_{i}$. The counting formula represents a counting random variable (CRV) whose range is the set of possible histograms that distribute $n=\left\lvert C_{\mathcal{L}}\right\rvert$ elements into $\prod_{i=1}^{k}\left\lvert range(B_{i})\right\rvert$ buckets. The state of $\gamma$ is the histogram function $h=\{(r_{i},n_{i})\}_{i=1}^{r}$ stating for each $r_{i}\in\times_{i=1}^{k}range(P_{i})$ the number $n_{i}$ of tuples $(P_{i})_{i}$ whose state is $r_{i}$.
If no restrictions apply, we omit $C$, and $n=\prod_{L\in\mathbf{L}}\left\lvert\mathcal{D}(L)\right\rvert$, where $\mathbf{L}$ is the set of logvars of the PRVs $B_{i}$. We do not define the operations to manipulate CRVs as we do not need them in this paper. The CRV corresponding to a clique gives us the number of occurrences for each possible instantiation of the PRVs in that clique. We give a small example for a CRV using an additional PRV $Friends(M,Y)$, with $M$ and $Y$ having the same domain, for illustrative purposes:
[…]
To have a PRV with two logvars, assume we have three PRVs, $Travel(M)$, $Friends(M,Y)$ and $Sick(Y)$ with $\left\lvert\mathcal{D(M)}\right\rvert=\left\lvert\mathcal{D}(Y)\right\rvert=8$. A possible state for the CRV #[Travel(M)\#[Travel(M), $Friends(M$, $Y)$, $Sick(Y)]$ is $\{(ttt,6)$, $(ttf,30)$, $(tft,4)$, $(tff,0)$, $(ftt,6)$, $(ftf,0)$, $(fft,0)$, (fff,18)}(fff,18)\} and we have $n=64$. The first histogram tells, that in this state, there are six tuples for which $Travel(a)=t$, $Friends(a,b)=t$, and $Sick(b)=t$ holds.
Example 4.1 illustrates that we need only eight buckets in a histogram regardless of the domain size of $M$ and $Y$. The number of buckets depends on the range values of the PRVs and not on the domain sizes of them. We now formalize the state representation, which is basically a set of CRVs per (maximal) clique in the relational cost graph:
[…]
We create one CRV $S_{i}$ for each (maximal) clique in the relational cost graph, counting the corresponding PRVs in the clique. For a single propositional RV in the relational cost graph, we use the RV for $S_{i}$ instead of a CRV. A state assigns a value to each $S_{i}$. The resulting state space is the set of possible states. We denote the state space by the tuple $(S_{i})_{i}$, which contains one CRV (or RV) per clique.
We give the representation of the state space for Example 3:
[…]
As the vertices are not connected in the relational cost graph, so the state representation is (#[Sick(M)](\#[Sick(M)], #[Travel(M)]\#[Travel(M)], $Epidemic)$.
The relational cost graph in Example 4.1 tells us that we can count the number of healthy persons separately from the number of travelling persons. In particular, we do not store which sick persons are travelling, since we do not need this information. We prove that our state representation is correct, i.e., retains the same semantics, namely exactly representing $\mathbf{S}$: {theorem} The representation in Definition 4.1 is correct.
Proof Sketch.
Given groundings for the state PRVs, we derive the histograms for the CRVs by counting the assignments for each parfactor. Given a representation as defined in Definition 4.1, we reconstruct, for each parfactor, the groundings of the PRVs by extracting the counts from the CRVs and instantiating the respective parfactors. We provide a full proof in Appendix A.4. ∎
To advance through an action to the next state, the action now has to use the same state representation, i.e., the action is specified on the counts for all PRVs of the current state for all parfactors the action is mentioned in:
[…]
The action $Restrict(M)$ uses the PRV $Travel(M)$. Thus, the mayor needs to specify how many persons of those (not) travelling are (not) allowed to travel. Suppose that in the current state five out of eight persons are travelling and two out of eight persons are sick. The action $Restrict(M)$ is defined over #[Travel(M),Restrict(M)]\#[Travel(M),Restrict(M)]. A concrete action is, e.g., $a=\{(tt,3)$, (ft,2)}(ft,2)\}, which means that three persons currently travelling are restricted from travelling and two persons not travelling. The action $a$ does not need to specify the counts of people no travel ban is imposed on ($tf$, ffff), as these are determined by $a$ and the current state.
The mayor no longer needs to specify individual persons (c.f. Example 3.1), but rather the number of persons (not) travelling, which are restricted from travelling. It is irrelevant on which exact persons the action is performed. With this action representation, we reduce the action space from exponential to polynomial, which we prove in Theorem 5 in the next section.
In the next subsection, we show how Foreplan uses the action space to compute the value function by solving a linear program.
4.2. Computing a Value Function
Let us have a look on how Foreplan computes the value function based on the introduced state representation. Foreplan uses the linear programming formulation given in Equation 2 to compute the value function. For the linear program, Foreplan uses the introduced state and action representations to iterate over all states and actions.
For instantiating the linear program in Equation 2, it is left to describe how Foreplan calculates the transition probability. Since we have full evidence provided, Foreplan evaluates each state CRV $S_{i}^{\prime}$ separately. For a fixed state CRV $S_{i}^{\prime}$, the value $s_{i}^{\prime}$ is fixed since the whole state space is iterated. For evaluating $P(s_{i}^{\prime}\mid s,a)$, Foreplan iterates over all possible assignment transitions, e.g., from each bucket in the histogram $s$ to each bucket in the histogram $s^{\prime}$. In our epidemic example, this is, e.g., the number of people getting sick and healthy. For each possible assignment transition, Foreplan calculates the transition probability: The probability for a representative transition is given by the involved parfactors, i.e., the parfactors defined over the PRVs mentioned in $S_{i}^{\prime}$. But this representative probability has to be weighted by how many times this assignment transition is applicable using the multinomial coefficient. Take for example the state in which five persons are sick. When three sick persons get healthy, this assignment transition has weight $5\choose 3$ as any three of the five persons can get healthy. The final probability of $P(s_{i}^{\prime}\mid s,a)=\sum_{t}w_{t}p_{t}$ for a fixed CRV is the sum over all possible assignment transitions $t$, each with representative probability $p_{t}$ and weight $w_{t}$. Since the state space is factored, the full transition probability $P(s^{\prime}\mid s,a)$ is factored as $P(s^{\prime}\mid s,a)=\prod_{i}P(s_{i}^{\prime}\mid s_{i},a)$. We provide a more detailed description and an example in Appendix A.5.
Let us have a closer look at the actual creation of Foreplan’s linear program: For each state and action combination, a constraint is added. Thus, Foreplan can take both, state and action, into account for generating only the necessary constraints. When an initial state is given, the reachable state space can be pruned, and Foreplan limits all computations to the pruned search space, that is why we are calling Foreplan a forward planner. Moreover, additional checks like mutual exclusion, capacities, or other constraints can be added, either on PRV level or on search space level. We make this more concrete with an example: Assume we have four persons travelling. Then, since the action uses the lifted state representation, the mayor can only restrict up to four persons, which currently travel, from travelling. The generation of constraints for the assumed state only iterates over the possible actions. Future work includes an analysis of what types of additional applicability checks are efficient implementable.
With Foreplan, we are able to cope with numerous indistinct objects and actions on collections of those objects. We do so by successfully applying lifting in the field of MDPs. While traditional approaches can represent actions on sets of objects, they fail to do so efficiently. Therein, the actions for each subset would be represented on their own, resulting in exponentially many actions. In the next section, we show the complexity of Foreplan.
5. Complexity Analysis of Foreplan
Having outlined Foreplan, we analyze the complexity of Foreplan in this section. We start by quantifying the state representation and using the complexity of the state representation to derive the runtime complexity of Foreplan.
We derive the following theorem about the size of the state representation from Definition 4.1. {theorem} The state representation is in $\mathcal{O}(c\times 2^{w})$.
Proof.
For each clique, the size of the histogram function is exponential in the number of vertices in the clique, as we enumerate all possible assignments. Thus, the size of the state representation is bounded by $c\times 2^{w}$. ∎
Note that $c$ and $w$ are independent of the domain sizes and determined only by the structure of the relational cost graph, and thus in general small. Also, $c$ is bounded by the number of PRVs in our parfactor model, as one parfactor is sufficient per PRV. Theorem 5 overapproximates the size of the state representation, as not all cliques have the same size and not both, $c$ and $w$ are large at the same time. Building on the size of the state representation, we give the complexity of the state and action space: {theorem} The state and action spaces are both polynomial in the number of objects and exponential in $c$ and $w$.
Proof.
We need to iterate over all possible instantiations of the state representation. For each clique (resp. CRV), the number of possible instantiations is polynomial in the number of objects and exponential in $w$. The joint state requires one instantiation per clique, resulting in the number $c$ of CRVs as exponent. The size of the action space is bounded by the size of the state space, as the action has to specify a (subset of a) state. ∎
Since Foreplan uses a linear program to compute the value function, we analyze the complexity of solving the linear program Foreplan builds. Linear programs can be solved in polynomial time w.r.t the variables and constraints 63499. Let us therefore take a closer look at the number of constraints and variables Foreplan generates:
{theorem}
The number of linear programming constraints and variables are polynomial in the state space.
Proof.
By Equation 2, Foreplan generates one variable per possible state and one constraint for each state and action combination. ∎
Plugging Theorem 5 in Theorem 5 leads to: {theorem} The runtime of Foreplan is polynomial in the number of objects and exponential in $c$ and $w$.
Therefore, with Foreplan, we have introduced an exact planner with a runtime polynomial instead of exponential w.r.t domain sizes for concurrent actions. The other parameters Foreplan is exponential in are mostly small and fixed. Thus, we have already achieved an exponential speed up.
Our main contribution Foreplan greatly advances the state of the art. However, as it is an exact algorithm, the value function is still not factorized. In case the runtime is of utter importance, we can also approxiamte the value function, following the idea of approx_factored_mdps, which has also been picked up by, e.g., 10.5555/3020336.3020398. By approximating the value function, we can prevent iterating over the joint state space leading to a blazingly fast, but approximate, version of Foreplan.
6. Foreplan: Even Faster by Approximation
While Foreplan runs in time polynomial in the number of objects, the runtime still depends exponentially on $c$. In this section, we present an approximation technique inspired by the Approximate Linear Programming (ALP) approx_factored_mdps approach to prevent iterating the whole state space. Our approach follows the same idea as ALP and ALP for first-order MDPs 10.5555/3020336.3020398. We first describe the approximation idea and then how Foreplan uses the approximation for our new representation and for concurrent actions. Last, we give bounds on the runtime and on the approximation quality. We call the approximate version Approximate Foreplan. Since Approximate Foreplan is an extension of our main contribution Foreplan, we refer to the Appendix for technical details.
Foreplan needs to iterate over the whole state space, because the value function maps from a state to the value of that state. We approximate the value function by a set $h_{i}$ of basis functions, whose scope is a subset of $\mathbf{S}$, $V\approx\sum_{i}w_{i}h_{i}$, where the goal is to find the most suitable weights $w_{i}$ approx_factored_mdps. Approximate Foreplan also needs the value of all possible next states in terms of the same approximation. Therefore, Approximate Foreplan makes use of the backprojections $g_{i}^{a}$ of the basis functions $h_{i}$ approx_factored_mdps, stating the influence of $x$ on the next state:
$g_{i}^{a}(x)=\sum_{x^{\prime}}P(x^{\prime}\mid x,a)\cdot h_{i}(x^{\prime})$ (4)
To compute the basis functions and backprojections lifted, we parameterize the basis functions analogue to the reward functions and calculate them lifted in the same way. The basis functions should capture the important dynamics in the model koller1999computing, most importantly the rewards 10.5555/3020336.3020398. As proposed by koller1999computing and picked up in 10.5555/3020336.3020398, we also use basis functions for capturing the rewards alongside a constant basis function:
[…]
The backprojections are computed lifted as defined next:
[…]
Let us apply the backprojection in our running example:
[…]
Suppose we have three sick persons and two healthy ones and are interested in the backprojection of $h_{1}$. Then, we have $G_{1}([(t,3),(f,2)],epi)=3\cdot g_{1}(true,epi)+2\cdot g_{1}(false,epi)$. We show the full calculation of all backprojections for our running example in Appendix B.3.
Approximate Foreplan precomputes all backprojections and then builds the following linear program approx_factored_mdps, whose result is the approximated value function and constitutes the result of Approximate Foreplan:
[…]
The $\alpha_{i}$’s are effectively coefficients for a linear combination over the $w_{i}$, stating how important the minimization of each $w_{i}$ is approx_factored_mdps; doi:10.1287/opre.51.6.850.24925. The maximum operator is not part of the definition of linear programs and is removed in an operation similar to variable elimination (VE) ve. In Appendix A.6, we provide an example for the removal procedure. Moreover, in Appendix B, we show a walkthrough of Approximate Foreplan on the epidemic example.
For the runtime analysis, we introduce the cost network briefly: The cost network for a constraint has a vertex for each appearing variable and there is an edge in the cost network between two vertices if the corresponding variables appear together in the same reward or basis function, or backprojection. For the complexity analysis, we use the structural graph parameter induced width, which provides an upper bound for the largest intermediate result dechter1999bucket:
{theorem}
Approximate Foreplan runs in time polynomial in the number of objects, polynomial in $c$ and exponential in the induced width of each cost network, when $w$ is bounded.
Proof Sketch.
The full proof is in Appendix A.7. Approximate Foreplan has to precompute the backprojections and solve the linear program in Equation 5. The first involves constantly many iterations of the state and action spaces. The second is linear in the action space and exponential in the induced width of each cost network approx_factored_mdps; dechter1999bucket. Because Approximate Foreplan does not iterate over the whole state space, but treats each clique independently in the maximum operator, the effective state space is no longer exponential in $c$, but polynomial, and the action space is bound by the state space. ∎
Most notably, $w$ and induced width in Theorem 6 are mostly small and fixed, leading to a polynomial runtime, as the growth in the number of objects is of more interest. In Appendix A.8, we combine the relational cost graph and the cost networks in a single total relational cost graph, and show that the runtime Approximate Foreplan is polynomial in the number of objects when the treewidth of the total relational cost graph is bounded.
Moreover, we can show that Approximate Foreplan and ALP compute the same solutions: {theorem} Given an rfMDP $R$, Approximate Foreplan and ALP are equivalent on $R$ and the grounded version of $R$, respectively.
Proof Sketch.
The full proof is in Appendix A.9. The objective function groups grounded basis functions and calculates them lifted when choosing appropiate $\alpha_{i}$. Each individual constraint is correct, because the lifted backprojections and lifted basis functions compute the same value as for the grounded model. The action representation covers the whole action space. ∎
With ALP and Approximate Foreplan being equivalent, we can transfer the approximation guarantee for ALP to Approximate Foreplan:
[…]
Approximate Foreplan provides the best approximation of the optimal value function in a weighted $\mathcal{L}_{1}$ sense, where the weights in the $\mathcal{L}_{1}$ norm are the state relevance weights $\alpha$.
Proof.
Since the claim holds for ALP approx_factored_mdps; doi:10.1287/opre.51.6.850.24925, the proof follows by the equivalence of ALP and Approximate Foreplan. ∎
With Approximate Foreplan, we reduce the runtime further from exponential in $c$ to polynomial in $c$. In the next section, we evaluate the runtimes of Foreplan and Approximate Foreplan empirically.
7. Empirical Evaluation
(Approximate) Foreplan runs in time polynomial in the number of objects, but other terms are unavoidably exponential. In contrast to current approaches, the exponential terms of both Foreplan variants depend only on the structure of the rfMDP and not on the number of objects. To underline our theoretical results, we evaluate (Approximate) Foreplan against ALP and an implementation of symbolic value iteration (VI) using extended algebraic decision diagrams (XADDs) pmlr-v162-jeong22a; taitler2022pyrddlgym for the epidemic example introduced in Example 3 as well as for the BoxWorld boutilier2001symbolic and a fully-connected SysAdmin approx_factored_mdps example, with the latter two extended to concurrent actions. We also assess the quality of the policy given by Approximate Foreplan. We use Python 3.12 and HiGHS for solving the linear programs highs. We run all implementations on an AMD EPYC 7452 32-Core Processor with 504 GB of RAM.
Figure 4 shows the runtime for the epidemic example for (Approximate) Foreplan, ALP and XADD Symbolic VI for up to 191 persons with a time limit of two hours. XADD Symbolic VI exceeds the time limit after ten persons, ALP after 15 persons. Foreplan times out after 20 persons and Approximate Foreplan after 191 persons. For ten persons, Foreplan is $91$ times faster than XADD Symbolic VI. Approximate Foreplan is even $14.000$ times faster, which are four orders of magnitude. For 15 persons, Foreplan is more than four times faster than ALP and Approximate Foreplan is more than $4.217$ times faster than ALP. For 20 persons, Approximate Foreplan is more than $4.077$ times faster than Foreplan. Thus, when using a symbolic solver, we can only solve the epidemic example for up to ten persons. With a factored and approximate approach, we can go up to 15 persons. In contrast, when using Foreplan, we can solve the epidemic example even for 20 persons and with Approximate Foreplan we can go further to 191 persons, which is ten times more than what ALP can solve.
We also run the four algorithms on a fully connected SysAdmin example with a timeout of two hours. XADD Symbolic VI times out after eleven computers and ALP after 12. With Foreplan, we can go up to 64 computers and with Approximate Foreplan even up to 94 computers. At eleven computers, Foreplan is more than $44.413$ times faster than XADD Symbolic VI and Approximate Foreplan is even more than $69.944$ times faster. At 12 computers, Foreplan is more than $8.359$ times faster than ALP and Approximate Foreplan even more than $16.173$ times. For 64 computers, Approximate Foreplan is more than $696$ times faster than Foreplan. In Appendix A.10, we present the runtime figure.
For the BoxWorld example, we use three cities and set a time limit of 15 hours. XADD Symbolic VI times out after two boxes and trains. ALP and Foreplan time out after nine boxes and trains. Approximate Foreplan manages to go up to 30 boxes and trains. At two boxes and trains, Foreplan is more than $83.227$ times faster than XADD Symbolic VI and Approximate Foreplan even more than $137.502$ times. At nine boxes and trains, Foreplan is more than three times faster than ALP and Approximate Foreplan is even more than $33.136$ times faster than ALP. In Appendix A.10, we present the runtime figure.
For assessing the quality of the policy given by Approximate Foreplan, we calculate the ratio of wrong actions over the total number of actions in every ground state. For ten persons, ALP and Approximate Foreplan both deviate in $1.2$ % of all actions. For two to ten persons, the deviation was $2.98$ % at most, with reducing errors when increasing the number of individuals. In all cases, ALP and Approximate Foreplan have an identical error. For SysAdmin, ALP and Approximate Foreplan both return the optimal policy in the tests going up to nine computers. For BoxWorld, we refer to Appendix A.10.
Overall, (Approximate) Foreplan achieves a speedup of several orders of magnitude and is able to compute policies for significantly more objects within the same time and memory limits. Moreover, the policy given by Approximate Foreplan is for more than $97$ % of all actions in the epidemic example correct and for all actions in the SysAdmin example.
Figure 4. Runtime of (Approximate) Foreplan, ALP and XADD Symbolic VI on the epidemic example for up to 191 persons with a time limit of two hours.
[…]
Graphical runtime comparison of Foreplan, Approximate Foreplan, Approximate Linear Programming and XADD Symbolic VI.
8. Conclusion
Propositional planning approaches struggle with having numerous indistinct objects and actions on subsets of those objects. While first-order MDPs can cope with numerous objects, current approaches for solving planning in them neglect concurrent actions and still need to represent actions for each subset individually, resulting in exponentially many actions. In this paper, we present Foreplan, a relational forward planner solving the exponential explosion by lifting the objects: Foreplan groups indistinguishable objects by a representative and carries out calculations on a representative-level. Afterwards, the result is projected to the whole group. Using histograms and focusing only on the number of objects an action is applied to, we effectively reduce the action space from exponential to polynomial in the number of objects.
In future work, we aim to develop a hybrid approach combining Foreplan and Golog levesque1997golog. With the forward search in Foreplan, we can identify states reachable from the initial state while the backwards search in Golog computes the exact optimal policy. Furthermore, the techniques from Foreplan can be transferred to first-order partially observable MDPs williams2005factored.
{acks}
The research for this paper was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2176 ’Understanding Written Artefacts: Material, Interaction and Transmission in Manuscript Cultures’, project no. 390893796. The research was conducted within the scope of the Centre for the Study of Manuscript Cultures (CSMC) at Universität Hamburg.
Appendix A Omitted Details
A.1. Examples for Hyperoperations in Relational Factored MDPs
Definition 3 requires the rewards to be represented as a sum of local reward functions. In fact, this requirement can be lifted to a broader one: Any operation $\oplus$ can run over a collection of local reward functions, as long as this operation has a hyperoperation $\bigoplus$, which performs $\oplus$ repeatedly. The most basic example is addition ($\oplus=+$) and multiplication as the corresponding hyperoperation ($\bigoplus=\cdot$). Moreover, these operations can be linked together in any way: We compute each operation on a collection of local reward functions lifted and then compute the overall result following the linking.
Let us first define some local reward functions and the reward function:
[…]
Assume we have PRVs $A(M)$ and $B(M)$, both with Boolean domain. We define the local reward functions $r_{1}(A(M))$ with $r_{1}(false)=2$ and $r_{1}(true)=5$. We further define $r_{2}(B(M))$ with $r_{2}(false)=-1$ and $r_{2}(true)=1$. We define the reward function as $\sum_{m\in M}r_{1}(m)\cdot\prod_{m\in M}r_{2}(m)$
The lifted computation of the reward function then calculates each basic term, i.e., $\sum_{m\in M}r_{1}(m)$ and $\prod_{m\in M}r_{2}(m)$, separately and combines the results using the linking operations, which is in this example the multiplication:
[…]
Assume $\left\lvert\mathcal{D}(M)\right\rvert=5$ and $A(M)$ is true for two of them and $B(M)$ is true for three of them. The lifted computation of the first term is then:
$\sum_{m\in M}r_{1}(m)=2\cdot r_{1}(true)+3\cdot r_{2}(false),$ (6)
and for the second term:
$\prod_{m\in M}r_{2}(M)=\left(r_{2}(true)\right)^{3}\cdot\left(r_{2}(false)\right)^{2}.$ (7)
The overall result is thus:
$\left(2\cdot r_{1}(true)+3\cdot r_{2}(false)\right)\cdot\left(\left(r_{2}(true)\right)^{3}\cdot\left(r_{2}(false)\right)^{2}\right).$ (8)
The hyperoperation is needed to reuse the result of one representative, lifting it to the group.
A.2. Formal Definition of Relational Cost Graph
To give a formal representation of the relational cost graph, we first define an undirected graph:
[…]
A graph $G=(V,E)$ is tuple of a set $V$ of vertices and $E\subseteq V\times V$ of edges. We require $E$ to be symmetric, i.e., $(a,b)\in E\implies(b,a)\in E$.
We further need some shorthand syntax to extract the logvars a PRV is defined over and to extract the PRVs a parfactor or local reward function is defined over:
[…]
Given a PRV $B$ defined over logvars $L_{1},\dots,L_{n}$, the notation $args(B)$ returns the logvars $L_{1},\dots,L_{n}$, i.e.,
$args(B)\coloneqq\{L_{1},\dots,L_{n}\}.$ (9)
We analogously define argsargs to return the arguments of a parfactor or local reward function: Given a parfactor $\phi$ defined over PRVs $B_{1},\dots B_{m}$, $args(\phi)$ returns those PRVs:
$args(\phi)\coloneqq\{B_{1},\dots,B_{m}\}.$ (10)
And given a local reward function $r$ defined over PRVs $A_{1},\dots,A_{k}$, $args(r)$ returns those PRVs:
$args(r)\coloneqq\{A_{1},\dots,A_{k}\}.$ (11)
Using these shorthand notations, we can give a formal definition of the relational cost graph:
[…]
Given an rfMDP with a set $\mathbf{R}$ of parameterized local reward functions, a set $\mathbf{\Phi}$ of factors, and a set $\mathbf{X}$ of PRVs, the relational cost graph is a graph $rcg=(V,E)$ with $V\coloneqq\mathbf{X}$ and the set $E$ of edges given by
$E\coloneqq\{(A,B)$ $\in\mathbf{X}\times\mathbf{X}\mid\left(args(A)\cap args(B)\neq\emptyset\right)$ (12)
$\land\left(\exists\phi\in\mathbf{\Phi}:a\in args(\phi)\land b\in args(\phi)\right)$ (13)
$\land\left(\exists r\in\mathbf{R}:a\in args(r)\land b\in args(r)\right)\}.$ (14)
A.3. Examples for the Relational Cost Graph
We illustrate the relational cost graph and later the state representation on a bit more complex setting:
[…]
Consider the PRVs $Sick(M)$ and $RemoteWork(M)$ and assume we have a parfactor defined over these two PRVs as well as the PRV $Sick^{\prime}(M)$ for the next state. Then, the relational cost graph consists of two vertices $Sick(M)$ and $RemoteWork(M)$ with an edge between them. Figure 5 shows the graph visually. Furthermore, we can understand why these two PRVs need to be counted together: A person is (not) sick in the next state dependent on that person being (not) sick and (not) working remote in the current state. We need both values for the correct transition probability.
Sick(M)Sick(M)$ReWo(M)$
Figure 5. Visual representation of the relational cost graph in Example A.3. We abbreviate $RemoteWork(M)$ by $ReWo(M)$.
[…]
The relational cost graph consists of two vertices Sick(M) and RemoteWork(M) with an edge between them.
We now apply the definition of a CRV to Example A.3.
[…]
The CRV for the clique consisting of $Sick(M)$ and $RemoteWork(M)$ has the structure #[Sick(M),RemoteWork(M)]\#[Sick(M),RemoteWork(M)] and a histogram for that CRV would consist of the four entries $n_{tt}$ for the number of people being sick and working remote, $n_{tf}$ for the number of people sick but not working remote, $n_{ft}$ for the number of people not sick but working remote, and $n_{ff}$ for the number of people not sick and not working remote.
We also provide an example for using our CRVs for counting PRVs defined over different parameters, where regular Counting Random Variables cannot be used:
[…]
Suppose we have a parfactor defined over PRVs $Sick(X),Friends(X,Y),Sick(Y)$. The respective CRV has structure #[Sick(X),Friends(X,Y),Sick(Y)]\#[Sick(X),Friends(X,Y),Sick(Y)] and we count for each (joint) assignment to the groundings of the PRVs how often this assignment occurs. Assume that $X$ has a domain with only one person $x$ and $Y$ with three, $y_{1}$ to $y_{3}$. Now further assume that $x$ and $y_{1}$ are sick and $x$ is friends with everyone except $y_{2}$. In other words, $x$ is friends with one sick and one healthy person and not friends with a sick person. Thus, the entries in the histogram for tttttt, ttfttf and tfttft are one, while all other ones are zero.
A.4. Proof of Correct State Representation
Theorem.
The representation in Definition 4.1 is correct.
Proof.
We first lay the foundation for the proof and then continue to show both directions of transformation.
By the definition of a parfactor model $G=\{g_{i}\}$, the full joint probability distribution is
$P_{G}$ $=\frac{1}{Z}\prod_{f\in gr(G)}f$ (15)
$=\frac{1}{Z}\prod_{g\in G}\prod_{f\in gr(g)}f,$ (16)
with $gr(g)$ referring to the groundings of a parfactor $g$. Without loss of generality, we fix a parfactor $g$. We split the PRVs $\mathcal{B}$ the factor $g$ is defined over into two sets: The set $\mathbf{B}_{in}$ of input PRVs, which contribute to the current state, and the set $\mathbf{B}_{out}$ of output PRVs, which contribute to the next state. We can ignore the set $\mathbf{B}_{out}$, because we only represent the current state and iterate over the next state later in Foreplan.
Given groundings for $g$, we count the occurrences of each assignment to the PRVs in $\mathbf{B}_{in}$. We split the counts into different CRVs according to Definition 4.1.
Given a state in CRVs, we have to instantiate all possible groundings respecting the current state representation. The PRVs in $\mathbf{B}_{in}$ are possibly split across different CRVs. We can combine any grounding of the PRVs in the different sets, as these PRVs do not share a parameter. ∎
A.5. Calculating Transition Probabilities
In this section, we show how Foreplan calculates the transition probabilities for the constraints in the linear program in Equation 2. For each state and action combination, Foreplan generates one constraint. Within this constraint, a sum is taken over all future states. We show how to calculate the required $P(s^{\prime}\mid s,a)$ for given state $s$, action $a$ and next state $s^{\prime}$. We assume that we have one parfactor per PRV in the next state.
Since the transition function is factored, Foreplan calculates the probability of each PRV in the next state separately. Thus, we only need to describe how Foreplan calculates $P(s_{j}^{\prime}\mid s,a)$ for a given current state, given action and given parfactor for state variable $s_{j}^{\prime}$ in the next state. In short, Foreplan first computes the state representation for the current and next state zoomed in at the given parfactor and second iterates over all possible transitions, summing their probabilities. We describe the two steps in more detail.
First, Foreplan computes the state representation only for the input PRVs (including the action PRV) of the parfactor from the given state representation and for the output PRV. Depending on the intertwinedness of the parfactors, the computation is just an extraction or summing out unneeded PRVs from the state representation. At the end of this step, Foreplan has a CRV for the input PRVs and another one for the output CRV for the parfactor.
Let us fix an order the buckets of the input CRV and denote the counts in each bucket by $k_{i},i=1,\dots,n$, where $n$ is the number of buckets.
In the second step, Foreplan iterates over all possible transitions, where one transition specifies transition counts $t_{i},i=1,\dots n$. The transition count $0\leq t_{i}\leq k_{i}$ specifies the number of objects transitioning from bucket $i$ to the true-assignment of the output CRV, $k_{i}-t_{i}$ to the false-assignment, respectively. Since the next state is given, the sum over the $t_{i}$ must be exactly the number of true-assignments in the output CRV. For a fixed transition, Foreplan calculates the probability $P_{t_{i},k_{i}}$ of a fixed bucket $i$ by
$P_{t_{i},k_{i}}=\binom{k_{i}}{t_{i}}\cdot\phi_{i\to true}^{t_{i}}\cdot\phi_{i\to false}^{k_{i}-t_{i}},$ (17)
where $\phi_{i\to true}$ denotes the probability for transitioning from bucket $i$ to truetrue and $\phi_{i\to false}$ for the transition to falsefalse. Both probabilites can be looked-up in the parfactor. For a parfactor for the output PRV $s_{j}$, the computation is thus
$P(s_{j}\mid s,a)=\sum_{(t_{i})_{i}}\prod_{i}P_{t_{i},k_{i}},$ (18)
where the sum $\sum_{(t_{i})_{i}}$ goes over all possible transitions.
We illustrate Equation 18 with our running example:
[…]
We show how to calculate the probability of the PRV $Travel$. Let us denote the number of persons travelling in the current state by $x$, the number of persons travelling in the next state by $x^{\prime}$, the number of persons travelling and restricted from travelling by $a_{1}$ and the number of persons not travelling and restricted from travelling by $a_{2}$. Then, the sum in Equation 18 goes over all $t_{1},t_{2},t_{3},t_{4}\geq 0$ with
$t_{1}+t_{2}+t_{3}+t_{4}$ $=x^{\prime},$ (19)
$t_{1}$ $\leq a_{1},$ (20)
$t_{2}$ $\leq x-a_{1},$ (21)
$t_{3}$ $\leq a_{2},$ (22)
$t_{4}$ $\leq m-x-a_{2},$ (23)
where $m$ is the number of persons in our example. We denote by $\phi(travel^{\prime},travel,restrict)$ the probability of a person travelling in the next state given that person is currently (not) travelling and (not) being restricted from travelling. Then, the values $P_{t_{i},k_{i}}$ are calculated by
$P_{t_{1},a_{1}}$ $=\binom{a_{1}}{t_{1}}\cdot\phi(t,t,t)^{t_{1}}\cdot\phi(f,t,t)^{a_{1}-t_{1}}$ (24)
$P_{t_{2},x-a_{1}}$ $=\binom{x-a_{1}}{t_{2}}\cdot\phi(t,t,f)^{t_{2}}\cdot\phi(f,t,f)^{x-a_{1}-t_{2}}$ (25)
$P_{t_{3},a_{2}}$ $=\binom{a_{2}}{t_{3}}\cdot\phi(t,f,t)^{t_{3}}\cdot\phi(f,f,t)^{a_{2}-t_{3}}$ (26)
$P_{t_{4},m-x-a_{2}}$ $=\binom{m-x-a_{2}}{t_{4}}\cdot\phi(t,f,f)^{t_{4}}$
$\quad\quad\cdot\phi(f,f,f)^{m-x-a_{2}-t_{4}}.$ (27)
A.6. Example of Removing a Maximum Operator
Let us briefly illustrate the step of removing the maximum operator. The removal is a two-phase process. In a first phase, ALP eliminates the variables and in a second phase, ALP generates the constraints for the linear programm along the elimination sequence. Suppose we have the function
$F=\max_{x_{1},x_{2}}f_{1}(x_{1})+f_{2}(x_{1},x_{2})+f_{3}(x_{2})$ (28)
in a linear program, e.g., $a\geq F$ or $a=F$, where $a\in\mathbb{R}$ or $a$ as a linear program variable. We start with the first phase. ALP eliminates $x_{1}$ by replacing $f_{1}$ and $f_{2}$ by a new function
$e_{1}(x_{2})=\max_{x_{1}}f_{1}(x_{1})+f_{2}(x_{1},x_{2}).$ (29)
ALP eliminates $x_{2}$ by replacing $e_{1}$ and $f_{3}$ by a new function
$e_{2}=\max_{x_{2}}e_{1}(x_{2})+f_{3}(x_{2}).$ (30)
Note that $e_{2}$ has an empty scope and evaluates therefore to a number. We continue with the second phase, in which ALP translates the elimination sequence into linear program constraints: In the linear program, ALP adds helping variables and constraints to enforce the maxima in the different terms approx_factored_mdps. For each function $e$ with domain $Z$, ALP adds the variables $u_{z}^{e}$ for each assignment $z$ to $Z$. The variable $u_{z}^{e}$ is supposed to yield the value of $e(z)$. For the initial functions $f_{i}$, in our case $f_{1}$, $f_{2}$, $f_{3}$, ALP simply adds $u_{z}^{f_{i}}=f_{i}(z)$ to the constraints of the linear program. Suppose we got the function $e(z)=\max_{x}e_{i}$ when eliminating some variable $x$. ALP then adds the constraints
$u_{z}^{e}\geq\sum_{i}u_{(z,x)}^{e_{i}}\ \forall z,x.$ (31)
For $e_{2}$, the generated constraints would be
$u^{e_{2}}\geq u_{x_{2}}^{e_{1}}+u_{x_{2}}^{f_{3}}$ (32)
for all possible values of $x_{2}$. We are interested in keeping the number of added constraints small, which is the aim of VE.
A.7. Proof of Runtime Complexity of Approximate Foreplan
We present the full proof for the runtime complexity of Approximate Foreplan:
Theorem.
Approximate Foreplan runs in time polynomial in the number of objects, polynomial in $c$ and exponential in the induced width of each cost network, when $w$ is bound.
Proof.
The runtime of Approximate Foreplan is the runtime of precomputing the backprojections and solving the linear program in Equation 5.
The precomputation involves to calculate the backprojections for each state and action combination. The backprojections are computed as given in Definition 6, which involves another iteration over the state space. Thus, the complexity of precomputing the backprojections is bounded by the state complexity raised to the power of three. However, the size of the state space is, in contrast to Foreplan, no longer exponential in $c$, but only polynomial in $c$: Because of the basis functions, the CRVs of the state representation are not iterated jointly, but independently. For the first part of Approximate Foreplan, we are left with a runtime complexity polynomial in the state space, which is polynomial in the number of objects and $c$, and exponential in $w$.
The second part is solving the linear program in Equation 5. Linear programs can be solved in time polynomial w.r.t the number of variables and constraints 63499. Therefore, we count the number of variables and constraints. Initially, we have one variable per basis function and one constraint per action. Then, the maximum operator in each action constraint is removed. The removal process follows the algorithm by approx_factored_mdps, who have shown their results for arbitrary functions and is therefore applicable. The removal process takes time exponential in the induced width of each cost network approx_factored_mdps; dechter1999bucket. Therefore, the runtime complexity for solving the linear program is polynomial in the number of actions and exponential in the induced width of each cost network.
The overall runtime complexity is (i) polynomial in $c$ and the number of objects, (ii) exponential in the $w$ and the induced width of each cost network. With $w$ bounded, the claim holds. ∎
A.8. More Runtime Theorems for Approximate Foreplan
We first define the total relational cost graph:
[…]
The total relational cost graph for a solution environment for an rfMDP contains a vertex for each (P)RV. Two vertices are connected by an edge if they occur together in a function or parfactor.
By definition, the total relational cost graph is a supergraph for all graphs of interest for the runtime complexity: {theorem} The following graphs are each subgraphs of the total relational cost graph:
  1. (1)
    relational cost graph
  2. (2)
    the cost network for each maximum constraint in Approximate Foreplan
Proof.
For both cases, we show that the total relational cost graph contains more vertices and edges than the respective definition requires. Thus, we can remove the superfluos vertices and edges to arrive at the respective subgraph.
We start with the relational cost graph. By Definition 4.1, the relational cost graph has a vertex for each PRV and an edge between two PRVs if they share a logvar and occur together in a parfactor, a parameterized local reward function, or a basis function. In particular, all these vertices and edges are introduced in the total relational cost graph too, when we connect two vertices once the corresponding PRVs occur together in a parfactor. We may add more edges than for the relational cost graph as we ignore the share a logvar condition.
We continue with the cost network for each maximum constraint. Let us take an arbitrary maximum constraint. The cost network consists of vertices for each variable appearing in the constraint and connects two vertices with an edge if the corresponding variables appear together in the same function. Thus, the total relational cost graph contains these edges too, as the respective variables occur together in a function. ∎
Since the induced width is the same as the treewidth minus one, we can give some more bounds: {theorem} If the total relational cost graph for an rfMDP has bounded treewidth, Approximate Foreplan runs in time polynomial in the number of objects of the rfMDP.
Proof.
By Theorem A.8, the relational cost graph is a subgraph of the total relational cost graph. As a subgraph can have the treewidth of the supergraph at most db122a323d0740a3b2b1580529f310e7, the treewidth of the relational cost graph is bounded. And if a graph has bounded treewidth, it also has bounded clique number, which is our $w$ db122a323d0740a3b2b1580529f310e7. Thus, Theorem 6 follows, leaving only the induced width of each cost network.
By Theorem A.8, each cost network is a subgraph of the total relational cost graph. The treewidth of each cost network is bounded by the treewidth of the total relational cost graph. As the induced width equals the treewidth minus one, it is bound, leaving no variable with exponential influence. ∎
A.9. Proof of Approximation Guarantee
We now provide the proof that Approximate Foreplan and ALP yield the same results on rfMDPs:
Theorem.
Given an rfMDP $R$, Approximate Foreplan and ALP are equivalent on $R$ and the grounded version of $R$, respectively.
Proof.
We first prove that the basis functions and backprojections evaluate to the same terms. Then, we investigate the setup of the linear programs.
The lifted basis functions accumulate multiple grounded ones. Evaluating a lifted basis function yields the same result as summing the grounded basis functions. For the backprojections, the case is very similar. Evaluating the lifted backprojections as in Definition 6 yields the same result as summing all grounded backprojections, because the objects grouped through PRVs are indistinguishable and share the same transition probabilities. Since we sum over all backprojections and basis functions in the linear program, the equivalence of the sums suffices.
For the linear program, we start with the objective function and continue with the constraints. The objective function used in Approximate Foreplan groups multiple grounded basis functions, which are used in ALP, together in one lifted basis function. Therefore, if we denote the weights in ALP by $\alpha^{\prime}_{i}$, we have the weights $\alpha_{i}=n_{i}\cdot\alpha^{\prime}_{i}$ for Foreplan, where $n_{i}$ stands for the number of grouped basis functions. Since the grouped basis functions are indistinguishable, the weights used in Approximate Foreplan are evenly distributed in ALP. Next, we have the constraints. Since the backprojections and basis functions in Approximate Foreplan evaluate to the same terms as the grounded functions in ALP, each individual constraint is correct. Furthermore, by Theorem 3, Approximate Foreplan covers the whole action space. ∎
A.10. Evaluation
Figure 6 shows the runtime of (Approximate) Foreplan, ALP and XADD Symbolic Value Iteration (VI) on the fully-connected SysAdmin approx_factored_mdps example with a time limit of two hours. Both versions of Foreplan clearly surpass ALP and XADD Symbolic VI.
For the BoxWorld boutilier2001symbolic example, we use trains instead of trucks and increase the number of boxes and trains evenly and simultaneously. Figure 7 shows the runtime of (Approximate) Foreplan, ALP and XADD Symbolic VI on the BoxWorld example with a time limit of 15 hours. Foreplan is more than $83.227$ times faster than the exact XADD Symbolic VI, which times out after two boxes and trains. Foreplan is at nine boxes and trains even more than three times faster than ALP. For approximation, Approximate Foreplan is more than $2.057$ times faster than ALP at nine boxes and trains. Moreover, Approximate Foreplan manages to go up to 30 boxes and trains within the same time limit.
Figure 6. Runtime (logscale) of (Approximate) Foreplan, ALP and XADD Symbolic VI on the SysAdmin example with a time limit of two hours. Only runs within these limits are shown. For Approximate Foreplan, we have chosen a bigger step size when no other algorithm was left for the number of computers.
[…]
The figure shows the runtime results of Foreplan, Approximate Foreplan, ALP and XADD Symbolic VI on the SysAdmin example.
Figure 7. Runtime (logscale) of (Approximate) Foreplan, ALP and XADD Symbolic VI on the BoxWorld example with a time limit of 15 hours. Only runs within these limits are shown. For Approximate Foreplan, we have chosen a bigger step size when no other algorithm was left for the number of boxes and trains.
[…]
The figure shows the runtime results of Foreplan, Approximate Foreplan, ALP and XADD Symbolic VI on the BoxWorld example.
The approximation error in the BoxWorld example is more complex. With our chosen basis functions, ALP and Approximate Foreplan both deviate in up to $85$ % of all actions, when tested on up to eight boxes and trains. Still, Approximate Foreplan is not worse than ALP, since both report the same coefficients for the basis functions, empirically and theoretically by Theorem 6. The reason lies in the reward and basis functions: The reward functions award points only for boxes in cities. Because we have chosen the basis functions identical to the reward functions and an additional constant basis function, the basis functions do not take boxes on trains into account, where the boxes may be unloaded in a more profitable city. Therefore, with better basis functions, we should be able to capture better solutions. Moreover, we highlight that Approximate Foreplan returns the same policy as ALP and is more than $30.000$ times faster.
Appendix B Walkthrough of Approximate Foreplan
We describe in this section how we can solve Example 3 with Approximate Foreplan. We first model the example formally and find the state representation. Afterwards, we precompute the backprojections and instantiate the linear program.
B.1. Modeling the Small Town
We model the setting in Example 3. The town population consists of three people. We refer to Figure 2 for the transition model. We define the parfactors $f_{1}$ and $f_{2}$ according to Tables 2 and 3. The parfactor $f_{3}$ sets the probability of an epidemic in the next state to $\frac{s+1}{5}$, where $s$ is the number of sick persons in the current state.
Table 2. Transition probabilities for each person giving the probability of travelling in the next timestep given whether the person is currently travelling and being restricted.
$Travel(M)$ $Restrict(M)$ $Travel^{\prime}(M)$
0 0 0.2
0 1 0.1
1 0 0.9
1 1 0.5
Table 3. Transition probabilities for each person giving the probability of being sick in the next timestep given whether the person is currently sick and there is an epidemic.
$Sick(M)$ EpidemicEpidemic $Sick^{\prime}(M)$
0 0 0.2
0 1 0.8
1 0 0.4
1 1 0.6
The mayor can restrict an arbitrary subset of the towns population from travelling. The local reward functions are $R_{1}(Sick(M))$ and $R_{2}(Travel(M))$, each for each person of the towns population. The function $R_{1}(Sick(M))$ evaluates to $1$ if the person is not sick and to 1-1 if the person is sick (c.f. Example 3.2):
$R_{1}(Sick(M))=\begin{cases}-1,&Sick(M)\\ 1,&\lnot Sick(M)\end{cases}$ (33)
The function $R_{2}(Travel(M))$ evaluates to $2$ if the person is travelling and to 0 otherwise:
$R_{2}(Travel(M))=\begin{cases}2,&Travel(M)\\ 0,&\lnot Travel(M)\end{cases}$ (34)
We set the discount factor $\gamma=0.9$. For choosing the basis functions, we go with the ones in Example 6, that is, $h_{0}=1$, $h_{1}=R_{1}$ and $h_{2}=R_{2}$.
B.2. State Space Representation
For using Approximate Foreplan, we first build the relational cost graph to find the state space representation. Afterwards, we can modify the actions to be compatible with the state space representation. The relational cost graph, depicted in Figure 3, contains two vertices, $Sick(M)$ and $Travel(M)$, because these are the only PRVs in the example. The relational cost graph does not contain any edge, because the two PRVs do not occur together in a parfactor, reward or basis function. Thus, the vertices $Sick(M)$ and $Travel(M)$ each are a clique of size one. Therefore, the state space representation contains two histograms along the value of the propositional random variable $Epidemic$. More formally, the state space representation is $([Sick(M)],[Travel(M)],Epidemic)$ (c.f. Example 4.1).
For the action $Restrict(M)$, the mayor no longer needs to specify on which person(s) she applies the travel ban(s) on, but has to define how many persons currently (not) travelling receive a travel restriction. Thus, the mayor needs to define the action histogram $[Travel(M),Restrict(M)]$ giving the numbers of people (not) travelling that are (not) restricted from travelling (c.f. Example 4.1).
B.3. Computing Backprojections
The first step of Foreplan is the precomputation of the backprojections. The backprojections are generally defined as
$g_{i}^{a}(x)=\sum_{x^{\prime}}P(x^{\prime}\mid x,a)\cdot h_{i}(x^{\prime}),$ (35)
where $x^{\prime}$ are the parameters of $h_{i}$, $x$ are the parents of $x$ in the transition model and $a$ is the selected action. For abbreviation, we use $S(M)$ for $Sick(M)$, $R(M)$ for $Restrict(M)$, $T(M)$ for $Travel(M)$ and $Epi$ for $Epidemic$ in this subsection. We start with the backprojection of $h_{0}=1$:
$g_{0}^{a}(x)=\sum_{x^{\prime}}P(x^{\prime}\mid x,a)\cdot h_{0}(x)=1\cdot\sum_{x^{\prime}}P(x^{\prime}\mid x,a)=1.$ (36)
Next, we backproject $h_{1}(Sick(M))$. Note that $Sick^{\prime}(M)$ is independent of the action in the transition model. Thus, the backprojection is independent of $a$:
$\begin{split}g_{1}(S(M),&Epi)=\\ &\sum_{x^{\prime}\in\{t,f\}}P(S^{\prime}(M)=x^{\prime}\mid S(M),epi)\cdot h_{1}(x^{\prime})\end{split}$ (37)
We plug the probabilities from Table 3 in and receive
$g_{1}(true,true)$ $=0.6\cdot(-1)+0.4\cdot 1$ $=$- $0.2,$ (38)
$g_{1}(true,false)$ $=0.4\cdot(-1)+0.6\cdot 1$ $=$ $0.2,$ (39)
$g_{1}(false,true)$ $=0.8\cdot(-1)+0.2\cdot 1$ $=$- $0.6,$ (40)
$g_{1}(false,false)$ $=0.2\cdot(-1)+0.8\cdot 1$ $=$ $0.6.$ (41)
The backprojection of $h_{2}$ is a bit more complex, as it includes the action. The general backprojection of $h_{2}$ is given by
$\begin{split}g_{2}^{R(M)}&(T(M))=\\ &\sum_{x^{\prime}\in\{t,f\}}P(T^{\prime}(M)=x^{\prime}\mid T(M),R(M))\cdot h_{2}(x^{\prime}).\end{split}$ (42)
We distinguish the two cases $R(M)=false$ and $R(M)=true$. We start with the first one, which leads to:
$g_{2}^{f}(t)$ $=\sum_{x^{\prime}\in\{t,f\}}P(T^{\prime}(M)=x^{\prime}\mid T(M)=t,R(M)=f)$ (43)
[…]
and
$g_{2}^{f}(f)$ $=\sum_{x^{\prime}\in\{t,f\}}P(t^{\prime}(P)=x^{\prime}\mid T(M)=f,R(M)=f)$ (44)
[…]
We continue with $R(M)=true$, leading to
$g_{2}^{t}(t)$ $=\sum_{x^{\prime}\in\{t,f\}}P(T^{\prime}(M)=x^{\prime}\mid T(M)=t,R(M)=t)$ (45)
[…]
$=1$
and
$g_{2}^{t}(f)$ $=\sum_{x^{\prime}\in\{t,f\}}P(T^{\prime}(M)=x^{\prime}\mid T(M)=f,R(M)=t)$ (46)
[…]
B.4. Instantiation of Linear Program
The second step of Approximate Foreplan is to instantiate the linear program. The linear program has three variables $w_{1}$, $w_{2}$ and $w_{3}$. The objective function is to minimize $\sum_{i=1}^{3}\alpha_{i}w_{i}$, with $\alpha_{i}$ being the relevance weights. We have one constraint for each possible action $a$:
[…]
where $x_{1},x_{2}\in\{0,1,2,3\}$ specify the number of persons being sick and travelling, respectively. The variable $x_{3}$ is Boolean and stores the truth value of $Epidemic$. With $G_{i}$ and $H_{i}$ we denote the lifted computation of $g_{i}$ and $h_{i}$. In the following, we write the lifted computation in terms of $x_{i}$ and $g_{i}$ or $h_{i}$. In the remainder of this subsection, we show the complete constraint generation for the linear program for one example action. We omit all other actions because of limited insights compared to one example instantiation and constraint generation. We show the constraint generation for the action of restricting nobody. We start with writing the constraint tailored to this setting and directly computing the value of basis functions and their backprojections lifted:
[…]
We continue with introducing functions to save steps later in variable elimination and constraint generation. We define four functions with different parameters, each collecting the respective terms:
$0\geq\max_{x_{1},x_{2},x_{3}}\{f_{1}(x_{1})+f_{2}(x_{2})+f_{3}(x_{1},x_{3})+f_{4}\},$ (49)
with
$f_{1}(x_{1})$ $=3-2x_{1}-w_{1}x_{1}h_{1}(t)-w_{1}\cdot(3-x_{1})\cdot h_{1}(f)$ (50)
$f_{2}(x_{2})$ $=2x_{2}+w_{2}\cdot(x_{2}\cdot 0.9\cdot g_{2}^{f}(t)-x_{2}h_{2}(t))$ (51)
$\quad\quad+w_{2}\cdot((3-x_{2})\cdot 0.9\cdot g_{2}^{f}(f))$
$f_{3}(x_{1},x_{3})$ $=0.9w_{1}\cdot(x_{1}g_{1}(t,x_{3})+(3-x_{1})\cdot g_{1}(f,x_{3}))$ (52)
$f_{4}$ $=-0.1w_{0}$ (53)
The only task we are left with is to remove the maximum operator. For that, we describe the two phases, variable elimination and constraint generation, in two different subsections.
B.4.1. Variable Elimination
We first eliminate $x_{2}$, leading to
$0\geq\max_{x_{1},x_{3}}\{f_{1}(x_{1})+e_{1}+f_{3}(x_{1},x_{3})+f_{4}\},$ (54)
with
$e_{1}=\max_{x_{2}}f_{2}(x_{2}).$ (55)
Next, we eliminate $x_{1}$:
$0\geq\max_{x_{3}}\{e_{2}(x_{3})+e_{1}+f_{4}\},$ (56)
with
$e_{2}(x_{3})=\max_{x_{1}}f_{1}(x_{1})+f_{3}(x_{1},x_{3}).$ (57)
Last, we eliminate $x_{3}$:
$0\geq e_{1}+e_{3}+f_{4},$ (58)
with
$e_{3}=\max_{x_{3}}e_{2}(x_{3}).$ (59)
B.4.2. Constraint Generation
We generate the constraints along the elimination order of the previous subsection. All the constraints listed in this subsection together replace the single constraint in Equation 47 for our example action of restricting nobody. We start with the constraints for the functions $f_{i}$ containing $w_{i}$s and continue with the functions $e_{i}$ obtained by eliminating variables:
$f_{1}(x_{1})$
$u_{0}^{f_{1}}$ $=3-3w_{1}$ (60)
$u_{1}^{f_{1}}$ $=3-2+w_{1}-2w_{1}=1-w_{1}$ (61)
$u_{2}^{f_{1}}$ $=3-4+2w_{1}-w_{1}=-1+w_{1}$ (62)
$u_{3}^{f_{1}}$ $=3-6+3w_{1}=-3+3w_{1}$ (63)
$f_{2}(x_{2})$
$u_{0}^{f_{2}}$ $=w_{2}\cdot(3\cdot 0.9\cdot 0.4)=3.1w_{2}$ (64)
$u_{1}^{f_{2}}$ $=2+w_{2}\cdot(1\cdot 0.9\cdot 1.8-1.2)+w_{2}\cdot(2\cdot 0.9\cdot 0.4)$
$=2+0.34w_{2}$ (65)
$u_{2}^{f_{2}}$ $=4+w_{2}\cdot(2\cdot 0.9\cdot 1.8-2\cdot 2)+w_{2}\cdot(1\cdot 1.9\cdot 0.4)$
$=4-0.04w_{2}$ (66)
$u_{3}^{f_{2}}$ $=6+w_{2}\cdot(3-0.9\cdot 1.8-3\cdot 2)=6-1.14w_{2}$ (67)
$f_{3}(x_{1},x_{3})$
$u_{00}^{f_{3}}$ $=w_{1}\cdot 3\cdot 0.9\cdot 0.6=1.62w_{1}$ (68)
$u_{01}^{f_{3}}$ $=-1.62w_{1}$ (69)
$u_{10}^{f_{3}}$ $=w_{1}\cdot 1\cdot 0.9\cdot 0.2+w_{1}\cdot 2\cdot 0.9\cdot 0.6=1.26w_{1}$ (70)
$u_{11}^{f_{3}}$ $=-1.26w_{1}$ (71)
$u_{20}^{f_{3}}$ $=w_{1}\cdot 2\cdot 0.9\cdot 0.2+w_{1}\cdot 1\cdot 0.9\cdot 0.6=0.9w_{1}$ (72)
$u_{21}^{f_{3}}$ $=-0.9w_{1}$ (73)
$u_{30}^{f_{3}}$ $=w_{1}\cdot 3\cdot 0.9\cdot 0.2=0.54w_{1}$ (74)
$u_{31}^{f_{3}}$ $=-0.54w_{1}$ (75)
$f_{4}$
$u^{f_{4}}=-0.1w_{0}$ (76)
$e_{1}$
$u^{e_{1}}$ $\geq u_{0}^{f_{2}}$ (77)
$u^{e_{1}}$ $\geq u_{1}^{f_{2}}$ (78)
$u^{e_{1}}$ $\geq u_{2}^{f_{2}}$ (79)
$u^{e_{1}}$ $\geq u_{3}^{f_{2}}$ (80)
$e_{2}(x_{3})$
$u_{0}^{e_{2}}$ $\geq u_{0}^{f_{1}}+u_{00}^{f_{3}}$ (81)
$u_{0}^{e_{2}}$ $\geq u_{1}^{f_{1}}+u_{10}^{f_{3}}$ (82)
$u_{0}^{e_{2}}$ $\geq u_{2}^{f_{1}}+u_{20}^{f_{3}}$ (83)
$u_{0}^{e_{2}}$ $\geq u_{3}^{f_{1}}+u_{30}^{f_{3}}$ (84)
$u_{1}^{e_{2}}$ $\geq u_{0}^{f_{1}}+u_{01}^{f_{3}}$ (85)
$u_{1}^{e_{2}}$ $\geq u_{1}^{f_{1}}+u_{11}^{f_{3}}$ (86)
$u_{1}^{e_{2}}$ $\geq u_{2}^{f_{1}}+u_{21}^{f_{3}}$ (87)
$u_{1}^{e_{2}}$ $\geq u_{3}^{f_{1}}+u_{31}^{f_{3}}$ (88)
$e_{3}$
$u^{e_{3}}$ $\geq u_{0}^{e_{2}}$ (89)
$u^{e_{3}}$ $\geq u_{1}^{e_{2}}$ (90)
Final Constraint
In the end, we add the final constraint
$0\geq u^{e_{1}}+u^{e_{3}}+u^{f_{4}}$ (91)
due to the introduction of our helper functions $f_{i}$.
B.4.3. Solving the Linear Program
After having generated all constraints for all actions, the linear program is fed into a solver to compute a solution. The template for the linear program looks like this:
[…]